OpenStudy (anonymous):

Solve the equation x^3-4x^2-x+4=0 and list all the solutions.

OpenStudy (anonymous):

\[x^3-4x^2-x+4=0\]

OpenStudy (anonymous):

what is the common factor in 2 first terms?

OpenStudy (anonymous):

x

OpenStudy (anonymous):

cant we take \(x^2\) as common factor?

OpenStudy (anonymous):

I guess so

OpenStudy (anonymous):

well start by factoring out \(x^2\) in 2 first term

OpenStudy (anonymous):

\[x^2(x-4)\]

OpenStudy (anonymous):

right

OpenStudy (anonymous):

now note that \(-x+4=-(x-4)\) and u have \[x^3-4x^2-x+4=x^2(x-4)-(x-4)\] what to do next?

OpenStudy (anonymous):

I don't know.

OpenStudy (anonymous):

we can take \(x-4\) as common factor...

OpenStudy (anonymous):

x-4(1)-(1)

OpenStudy (anonymous):

(x-4)(x^2-1)=0 x-4=0 or x2-1=0 x=4 & x=1 & x=-1

OpenStudy (anonymous):

Thanks